Torre Archimede, room 427 Via Trieste 63 35121 Padua Italy

November 3, 2011

Prof. Dr. Annie Cuyt Universiteit Antwerpen Department of Mathematics and Computer Science Campus Middelheim, Building G Middelheimlaan 1 B-2020 Antwerpen Belgium

Dear Professor

I am writing to you on behalf of the Organizing Committee of the 3rd Dolomites Workshop on Constructive Approximation and Application, that will be held in Alba di Canazei (Trento, Italy), from September 9 to 14, 2012.

We are pleased to invite you to give a one-hour plenary talk at the workshop and we would be honored if you will be able to accept. With regard to expenses, we will be able to reimburse you up to an amount of 900 Euros. For your information, we mention that the cost of a nice hotel, including half pension (breakfast and dinner), in Canazei during the period of the conference is typically about 70 Euros and we hope that the amount mentioned above is sufficient to cover all, or at the very least, most of your expenses.

If indeed you are able to come, we would be grateful if you could send us a preliminary title for your talk.

Best regards,

Alvise Sommariva